A student is running at a constant velocity of 5 m/s in an effort to catch a bus parked at a station. When he is 10 m away from the bus, the bus pulls away with constant acceleration of 1.2 m/s2. How much time will it take the student to catch the bus if he keeps running with the same velocity?


1
Expert's answer
2021-09-06T11:33:07-0400

Given:

"v=5\:\rm m/s"

"a=1.2\:\rm m/s^2"

"l=10\:\rm m"


The distance traveled by a student

"d_s=vt"

The distance traveled by a bus

"d_b=at^2\/2"

We get

"d_s+l=d_b"

"vt+l=at^2\/2"

"5t+10=0.6t^2"

"0.6t^2-5t-10=0"

Root

"t=3.3\:\rm s"